You are selling tickets for a musical at your local community college. Student tickets cost $5 and general admission tickets cost $8. If you sell 500 tickets and collect $3475, how many student tickets and how many general admission
Answer by stanbon(75887)

Student tickets cost $5 and general admission tickets cost $8. If you sell 500 tickets and collect $3475, how many student tickets and how many general admission
----------------
Let # of student tickets sold be "x" ; Value of those is 5x dollars
Number of gen. admis. sold is "500-x"; Value of those is
8(500-x)=4000-8x dollars.
------------
EQUATION:
value + value = 3475 dollars
5x+4000-8x=3475
-3x=-525
x=175 (number of student tickets sold)
500-x=325 (number of gen adm. tickets sold)
Cheers,
Stan H.